OpenStudy (anonymous):

Use the distributive property law to factor each of the following. Check by multiplying. 1. 30r + 42s

ganeshie8 (ganeshie8):

hint : 30 = 6x5

ganeshie8 (ganeshie8):

hint2 : 42 = 6x7

OpenStudy (anonymous):

hint3: 6(5r+7s) :D

ganeshie8 (ganeshie8):

lol :)

OpenStudy (anonymous):

I still dont get it

ganeshie8 (ganeshie8):

use the first two hints, write 30 as 6x5, and 42 as 6x7

OpenStudy (anonymous):

okay. then what?

ganeshie8 (ganeshie8):

30r + 42s 6x5r + 6x7s

ganeshie8 (ganeshie8):

do you see 6 as common in both terms ?

ganeshie8 (ganeshie8):

pull out 6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then what?

ganeshie8 (ganeshie8):

30r + 42s 6x5r + 6x7s ^ ^

ganeshie8 (ganeshie8):

pull out and see wat u get :)

OpenStudy (anonymous):

He won't get it :D

OpenStudy (anonymous):

you get 5r+7s

OpenStudy (anonymous):

Zarax: I'm a girl...

ganeshie8 (ganeshie8):

yes ! where does 6 go... keep it also

OpenStudy (anonymous):

so it is 6(5r=7s) ?

ganeshie8 (ganeshie8):

Excellent !

ganeshie8 (ganeshie8):

hope you meant 6(5r+7s)

OpenStudy (anonymous):

haha yeah thats what I meant. Thanks so much! :)

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

I'm sorry. Factoring allows us to write expressions in an easy and understandable way. You get rid of the big numbers. instead of writing 300x + 600y you write 300(x+2y). Those two are the same. But in the second one you can understand it easily. So 30r+42s = 6(5r+7s) and 6(5r+7s) = 30r+42s